# Flexible Multitask Learning with Factorized Diffusion Policy

### v1

@@ Abstract
Multitask learning poses significant challenges due to the highly multimodal and diverse nature of robot action distributions. However, effectively fitting policies to these complex task distributions is often difficult, and existing monolithic models often underfit the action distribution and lack the flexibility required for efficient adaptation. We introduce a novel modular diffusion policy framework that factorizes complex action distributions into a composition of specialized diffusion models, each capturing a distinct sub-mode of the behavior space for a more effective overall policy. In addition, this modular structure enables flexible policy adaptation to new tasks by adding or fine-tuning components, which inherently mitigates catastrophic forgetting. Empirically, across both simulation and real-world robotic manipulation settings, we illustrate how our method consistently outperforms strong modular and monolithic baselines. Website: chaoqi-liu.com/factorpolicy.
@@ I Introduction
Imitation learning has emerged as a powerful paradigm for acquiring complex robotic manipulation skills [chi2024diffusionpolicy, wang2024poco]. However, extending this success to multitask settings remains a significant challenge. As the variety of tasks increases, the underlying action distribution becomes highly multimodal and diverse, often involving distinct control strategies across different objects. Traditional monolithic policies often struggle to generalize across tasks, represent multiple behavior modes, or adapt efficiently to new skills [wang2024poco, ha_scaling_2023, chang2020sight].
@@
To address these limitations, modular policy architectures, most notably Mixture-of-Experts (MoE) models [shazeer_outrageously_2017, lin2024moma], have emerged as a promising direction. By decomposing the policy into specialized components, modular methods improve scalability and reuse across tasks [wang2024poco, yang2020multi, wang2024sparse_dp_moe, reuss2024mode, chen2025multimodalmanipulationmultimodalpolicy, huang2025moe]. Yet, existing MoE-based approaches often suffer from training instability [lin2024moma], lack a principled probabilistic formulation, and produce expert modules with unclear or overlapping roles [wang2024sparse_dp_moe, du2024reducereuserecyclecompositional], limiting their interpretability.
@@
We propose Factorized Diffusion Policy (FDP), a simple yet effective modular policy architecture. FDP decomposes the policy into multiple diffusion components (Fig. 1a), each capturing a distinct behavioral mode, which are dynamically composed at inference time via an observation-conditioned router (Fig. 1c). Instead of discrete expert selection as in standard MoE architectures, FDP uses continuous score aggregation, enabling stable training, preventing routing imbalance, and promoting clearer specialization across components. FDP is grounded in compositional diffusion modeling [du2024reducereuserecyclecompositional, du2024compositional, liu2023compositionalvisualgenerationcomposable], where aggregating scores corresponds to sampling from the product of distributions, providing a principled probabilistic interpretation and a natural formulation as constraint satisfaction. The modular structure further enables efficient task adaptation: we extend the policy by introducing new diffusion components initialized via upcycling [lin2024moma] from existing components (Fig. 1b), allowing efficient skill expansion without retraining the entire policy. This factorization improves multitask learning and supports scalable adaptation.
@@
We validate FDP through extensive experiments in simulation benchmarks MetaWorld [yu2021metaworldbenchmark] and RLBench [james2019rlbenchbenchmark], and further demonstrate its practical benefits in real-world robotic manipulation. Our contributions are summarized as follows: (1) We introduce a modular diffusion policy architecture that composes specialized components via observation-conditioned compositional sampling. (2) We demonstrate that our compositional framework improves multitask performance and enables sub-skill decomposition across diffusion modules. (3) We propose a simple and effective strategy for adapting to new tasks by selectively tuning or augmenting existing components, achieving superior sample efficiency and modular reuse.
@@ II Related Works
Diffusion Models for Robotics. Diffusion models have emerged as a powerful tool for modeling complex distributions, achieving strong performance in image [ho2020ddpm, pmlr-v139-nichol21a, ramesh2022hierarchicaltextconditionalimagegeneration] and video generation [ho_video_2022, wang2025hma]. Their stable training and generative flexibility have led to increasing adoption in robotic domains, including video-conditioned policy learning [du_learning_2023, ajay_compositional], grasp synthesis [urainSEDiffusionFieldsLearning2023a], bimanual manipulation [chen2025bimanual], tool use [chen2025toolasinterface], trajectory planning [jannerPlanningDiffusionFlexible2022a, ajayConditionalGenerativeModeling2023, carvalhoMotionPlanningDiffusion2023], and closed-loop visuomotor control. Diffusion Policy (DP) [chi2024diffusionpolicy] demonstrated that diffusion models can be used to learn reactive visuomotor policies from demonstrations, achieving state-of-the-art performance in single-task imitation learning.
Multitask Imitation Learning and Adaptation. Traditional approaches to multitask imitation learning often rely on monolithic networks [kim_openvla_2024, octomodelteam2024octo] or language-conditioned policies [ha_scaling_2023, reussGoalConditionedImitationLearning2023a], which limit scalability, reusability, and interpretability. While early research established modular architectures to improve task decomposition [devin2016learningmodularneuralnetwork, andreas2017modularmultitaskreinforcementlearning], modern Sparse Diffusion Policy (SDP) [wang2024sparse_dp_moe] and variational distillation methods for MoE [zhou2024variationaldistillationdiffusionpolicies] extend this modular principle by introducing MoE layers in diffusion models, activating sparse expert sets based on observations. While this modular design enables expert reuse and policy expansion, it suffers from instability and load imbalance [lin2024moma]. Mixture-of-Denoising-Experts (MoDE) [reuss2024mode] conditions expert routing on noise level, distributing learning across noise levels, making its experts less interpretable or transferable across tasks. In contrast, FDP composes diffusion models through continuous score aggregation, avoiding hard expert selection and ensuring all components are jointly optimized. This promotes stable optimization, clear specialization, and better load balancing. While maintaining modular extensibility like MoE designs, FDP allows efficient adaptation by adding new components without overwriting prior skills.
@@ III-A Probabilistic Policy Modeling
$p(\mathbf{a}_{t}\mid\mathbf{o}_{t})\propto\prod_{i}p_{i}(\mathbf{a}_{t}\mid\mathbf{o}_{t})^{w_{t,i}},$
where $\{w_{t,i}\}$ are observation-dependent weights associated with each component distribution. Intuitively, $p(\mathbf{a}_{t}\mid\mathbf{o}_{t})$ represents the intersection (logical AND) of individual distributions, assigning high likelihood to samples commonly favored by all component distributions. Moreover, each diffusion component $p_{i}(\mathbf{a}_{t}\mid\mathbf{o}_{t})$ can be interpreted as imposing a behavioral constraint (e.g., collision avoidance, precise grasping) [yang2023compositionaldiffusionbasedcontinuousconstraint]. The composed distribution thus captures the intersection of constraints, naturally framing action generation as constraint satisfaction while maintaining a probabilistic interpretation.
Denoising Diffusion Probabilistic Model (DDPM) framework [ho2020ddpm] is adopted to model each component distribution $p_{i}(\mathbf{a}_{t}\mid\mathbf{o}_{t})$. To sample from each component, we start from a noisy action sample $\mathbf{a}^{K}_{t,i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and iteratively refine it using a noise prediction network $\bm{\varepsilon}_{\theta_{i}}(\mathbf{a}^{k}_{t,i},\mathbf{o}_{t},k)$, progressively denoising over $k$ steps:
@@
where $\alpha_{k}$, $\gamma_{k}$, and $\sigma_{k}$ define the noise schedule. This process closely resembles Stochastic Langevin Dynamics [welling2011bayesian], with $\bm{\varepsilon}_{\theta_{i}}$ estimating the score function $\nabla\log p_{i}(\mathbf{a}_{t,i}\mid\mathbf{o}_{t})$ [vincent2011score_matching].
@@
where $\mathbf{a}^{0}_{t,i}$ is a clean trajectory sample valid under distribution $p_{i}(\mathbf{a}_{t}\mid\mathbf{o}_{t})$. Minimizing this loss teaches the network to progressively denoise noisy actions conditioned on observations.
@@ III-B Compositional Sampling and Routing
We next discuss how can we sample from the actual action distribution $p(\mathbf{a}_{t}\mid\mathbf{o}_{t})$ given DDPM formulation of component distributions $\{p_{i}(\mathbf{a}_{t}\mid\mathbf{o}_{t})\}$, as well as how to automatically discover each component distribution and optimize corresponding diffusion models jointly.
One way of viewing the composition of distributions is through the lens of energy-based models (EBM) [du2019implicitEBM]. Assume weights $\{w_{t,i}\}$ are given, and each weighted component distribution is parameterized as $p_{i}(\mathbf{a}_{t}\mid\mathbf{o}_{t})\propto e^{-w_{t,i}E_{i}}$, then the actual action distribution can be expressed as $p(\mathbf{a}_{t}\mid\mathbf{o}_{t})\propto e^{-\sum_{i}w_{t,i}E_{i}}$ [du2019implicitEBM]. Therefore, iterative sampling can be performed via Langevin dynamics:
@@
where $\gamma_{k}$ controls the step size and $\xi_{k}$ introduces Gaussian noise. Note that we can bridge EBM with score-matching diffusion models [wang2024poco, du2024reducereuserecyclecompositional, liu2023compositionalvisualgenerationcomposable, su2024compositionalimagedecompositiondiffusion], which updates Equ. 2 as
@@
Compared to discrete MoE routing, our compositional approach avoids routing instability and expert imbalance [lin2024moma] by assigning continuous, observation-dependent weights to all components, rather than selecting a hard subset. In MoE, only a few experts are activated at each step, which can lead to underutilization of some experts and overfitting or saturation in others, especially when routing distributions are sharp or poorly calibrated. In contrast, our method aggregates contributions from all components via soft score-weighted composition, ensuring all modules remain active during optimization. Additionally, because all components participate in every training step, they receive gradient signals consistently, which encourages functional specialization.
@@ III-C Multitask Learning and Adaptation
Multitask Learning. This factorization is particularly well-suited for multitask imitation learning, where action distributions are inherently multimodal due to diverse object properties, contact dynamics, and task goals. In contrast to monolithic policies that must capture all modes simultaneously, FDP distributes complexity evenly across diffusion components, each modeling a coherent subspace of behaviors. Unlike MoE policies, where skills may span combinations of experts across layers, our formulation yields disentangled sub-skills.
Adapting to New Tasks. The modularity of FDP also enables efficient adaptation to unseen tasks. Instead of retraining the full model, we adapt by introducing a new diffusion component $\bm{\varepsilon}_{\theta_{\text{new}}}$, initialized via upcycling [lin2024moma] – copying weights from existing components. The updated score function becomes:
@@ IV-A Experiments Setup
((b)) Illustration of hang-X
((c)) Illustration of cube-X
@@
We evaluate policies on 30+ tasks in simulation across MetaWorld [yu2021metaworldbenchmark], RLBench [james2019rlbenchbenchmark], and LIBERO [liu2023libero]. Real-world experiments use a UR5e arm with a Robotiq gripper and a RealSense D415 camera (Fig. 2(a)). We evaluate policies on 4 distinct tasks: cube red, cube blue, hang low, and hang high. The tasks cube-X involve picking up a cube of color X from the tabletop and placing it into a designated bowl. The hang-X tasks require the robot to grasp a mug from the tabletop and precisely hang it on the X branch of a mug stand positioned on the table. Illustrations and setups of these real-world tasks are shown in Fig. 2. For MetaWorld and RLBench, goals are implicitly specified by the scene configuration, whereas for LIBERO and real-world experiments, tasks are identified using discrete integer task indices.
@@ IV-B Implementations
All policies take RGB images and joint angles as input and predict absolute joint angle trajectories. A history window of size 2 is used, with 16-step trajectories predicted and 8 steps executed. While we employ DDPM [ho2020ddpm], FDP is solver-agnostic; alternative samplers like DDIM [song2022ddim] offer comparable performance with reduced inference latency. We compare FDP against three baselines: DP [chi2024diffusionpolicy], a monolithic diffusion policy; SDP [wang2024sparse_dp_moe], a MoE-based diffusion policy with observation-conditioned routing; and MoDE [reuss2024mode], a MoE variant with routing based on noise levels. We follow the original configurations used in each baselines, and proportionally reduce the model size of MoDE to match others. We refer readers to the original papers for more details on architecture and training. In FDP, four U-Net diffusion modules are composed. For adaptation, we adopt the upcycling strategy [lin2024moma] to initialize new MoE experts or diffusion components from existing ones.
@@ IV-C Multitask Learning
Simulation. We evaluate FDP on 6 MetaWorld tasks (25 demonstrations each) and 8 RLBench tasks (50 demonstrations each). All methods are evaluated over 40 rollouts per task, with results shown in Table I. The DP baseline performs surprisingly well, particularly on tasks like drawer open, assembly, and hammer, which primarily involve reaching and grasping and exhibit fewer multimodal behaviors – making them easier to solve with a single model. Among modular baselines, SDP underperforms due to instability common in training MoE architectures [lin2024moma]: too few experts limit expressiveness, while too many can cause overfitting and noisy routing. MoDE performs reasonably by routing based on the noise level, but still inherits instability from MoE training [reuss2024mode]. In contrast, FDP’s compositional structure avoids abrupt routing decisions by continuously composing diffusion component outputs via score-weighted aggregation, which enables stable training and more balanced component specialization of the multimodal action distributions.
Policy Cube Red Hang Low Avg.
DP 0.700 0.800 0.750
SDP 0.750 0.650 0.700
MoDE 0.700 0.800 0.750
FDP 0.750 0.850 0.800

### v2

@@ Abstract
Multitask learning poses significant challenges due to the highly multimodal and diverse nature of robot action distributions. However, effectively fitting policies to these complex task distributions is often difficult, and existing monolithic models often underfit the action distribution and lack the flexibility required for efficient adaptation. We introduce a novel modular diffusion policy framework that factorizes complex action distributions into a composition of specialized diffusion models, each capturing a distinct sub-mode of the behavior space for a more effective overall policy. In addition, this modular structure enables flexible policy adaptation to new tasks by adding or fine-tuning components, which inherently mitigates catastrophic forgetting. Empirically, across both simulation and real-world robotic manipulation settings, we illustrate how our method consistently outperforms strong modular and monolithic baselines. Website: factorized-diffusion-policy.github.io.
@@ I Introduction
Imitation learning has emerged as a powerful paradigm for acquiring complex robotic manipulation skills [10, 37, høeg2025hybriddiffusionsimultaneoussymbolic, 8, 7]. However, extending this success to multitask settings remains a significant challenge. As the variety of tasks increases, the underlying action distribution becomes highly multimodal and diverse, often involving distinct control strategies across different objects. Traditional monolithic policies often struggle to generalize across tasks, represent multiple behavior modes, or adapt efficiently to new skills [37, 16, 6].
@@
To address these limitations, modular policy architectures, most notably Mixture-of-Experts (MoE) models [31, 22], have emerged as a promising direction. By decomposing the policy into specialized components, modular methods improve scalability and reuse across tasks [37, 41, 39, 30, 7]. Yet, existing MoE-based approaches often suffer from training instability [22], lack a principled probabilistic formulation, and produce expert modules with unclear or overlapping roles [39, 12], limiting their interpretability.
@@
We propose Factorized Diffusion Policy (FDP), a simple yet effective modular policy architecture. FDP decomposes the policy into multiple diffusion components (Fig. 1a), each capturing a distinct behavioral mode, which are dynamically composed at inference time via an observation-conditioned router (Fig. 1c). Instead of discrete expert selection as in standard MoE architectures, FDP uses continuous score aggregation, enabling stable training, preventing routing imbalance, and promoting clearer specialization across components. FDP is grounded in compositional diffusion modeling [12, 13, 25], where aggregating scores corresponds to sampling from the product of distributions, providing a principled probabilistic interpretation and a natural formulation as constraint satisfaction. The modular structure further enables efficient task adaptation: we extend the policy by introducing new diffusion components initialized via upcycling [22] from existing components (Fig. 1b), allowing efficient skill expansion without retraining the entire policy. This factorization improves multitask learning and supports scalable adaptation.
@@
We validate FDP through extensive experiments in simulation benchmarks MetaWorld [43] and RLBench [19], and further demonstrate its practical benefits in real-world robotic manipulation. Our contributions are summarized as follows: (1) We introduce a modular diffusion policy architecture that composes specialized components via observation-conditioned compositional sampling. (2) We demonstrate that our compositional framework improves multitask performance and enables sub-skill decomposition across diffusion modules. (3) We propose a simple and effective strategy for adapting to new tasks by selectively tuning or augmenting existing components, achieving superior sample efficiency and modular reuse.
@@ II Related Works
II-1 Diffusion Models for Robotics
Diffusion models have emerged as a powerful tool for modeling complex distributions, achieving strong performance in image [17, 26, 28] and video generation [18, 38]. Their stable training and generative flexibility have led to increasing adoption in robotic domains, including video-conditioned policy learning [15, 2], grasp synthesis [35], bimanual manipulation [8], tool use [9], trajectory planning [20, 1, 5], and closed-loop visuomotor control. Diffusion Policy (DP) [10] demonstrated that diffusion models can be used to learn reactive visuomotor policies from demonstrations, achieving state-of-the-art performance in single-task imitation learning.
II-2 Multitask Imitation Learning and Adaptation
Traditional approaches to multitask imitation learning often rely on monolithic networks [21, 34] or language-conditioned policies [16, 29], which limit scalability, reusability, and interpretability. While early research established modular architectures to improve task decomposition [11, 3], modern Sparse Diffusion Policy (SDP) [39] and variational distillation methods for MoE [44] extend this modular principle by introducing MoE layers in diffusion models, activating sparse expert sets based on observations. While this modular design enables expert reuse and policy expansion, it suffers from instability and load imbalance [22]. Mixture-of-Denoising-Experts (MoDE) [30] conditions expert routing on noise level, distributing learning across noise levels, making its experts less interpretable or transferable across tasks. In contrast, FDP composes diffusion models through continuous score aggregation, avoiding hard expert selection and ensuring all components are jointly optimized. This promotes stable optimization, clear specialization, and better load balancing. While maintaining modular extensibility like MoE designs, FDP allows efficient adaptation by adding new components without overwriting prior skills.
@@ III-A Probabilistic Policy Modeling
$p(\mathbf{a}_{t}|\mathbf{o}_{t})\propto\prod_{i}p_{i}(\mathbf{a}_{t}|\mathbf{o}_{t})^{w_{t,i}},$
where $\{w_{t,i}\}$ are observation-dependent weights associated with each component distribution. Intuitively, $p(\mathbf{a}_{t}|\mathbf{o}_{t})$ represents the intersection (logical AND) of individual distributions, assigning high likelihood to samples commonly favored by all component distributions. Moreover, each diffusion component $p_{i}(\mathbf{a}_{t}|\mathbf{o}_{t})$ can be interpreted as imposing a behavioral constraint (e.g., collision avoidance, precise grasping) [42]. The composed distribution thus captures the intersection of constraints, naturally framing action generation as constraint satisfaction while maintaining a probabilistic interpretation.
Denoising Diffusion Probabilistic Model (DDPM) framework [17] is adopted to model each component distribution $p_{i}(\mathbf{a}_{t}|\mathbf{o}_{t})$. To sample from each component, we start from a noisy action sample $\mathbf{a}^{K}_{t,i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and iteratively refine it using a noise prediction network $\bm{\varepsilon}_{\theta_{i}}(\mathbf{a}^{k}_{t,i},\mathbf{o}_{t},k)$, progressively denoising over $k$ steps:
@@
where $\alpha_{k}$, $\gamma_{k}$, and $\sigma_{k}$ define the noise schedule. This process closely resembles Stochastic Langevin Dynamics [40], with $\bm{\varepsilon}_{\theta_{i}}$ estimating the score function $\nabla\log p_{i}(\mathbf{a}_{t,i}|\mathbf{o}_{t})$ [36].
@@
where $\mathbf{a}^{0}_{t,i}$ is a clean trajectory sample valid under distribution $p_{i}(\mathbf{a}_{t}|\mathbf{o}_{t})$. Minimizing this loss teaches the network to progressively denoise noisy actions conditioned on observations.
@@ III-B Compositional Sampling and Routing
We next discuss how can we sample from the actual action distribution $p(\mathbf{a}_{t}|\mathbf{o}_{t})$ given DDPM formulation of component distributions $\{p_{i}(\mathbf{a}_{t}|\mathbf{o}_{t})\}$, as well as how to automatically discover each component distribution and optimize corresponding diffusion models jointly.
One way of viewing the composition of distributions is through the lens of energy-based models (EBM) [14]. Assume weights $\{w_{t,i}\}$ are given, and each weighted component distribution is parameterized as $p_{i}(\mathbf{a}_{t}|\mathbf{o}_{t})\propto e^{-w_{t,i}E_{i}}$, then the actual action distribution can be expressed as $p(\mathbf{a}_{t}|\mathbf{o}_{t})\propto e^{-\sum_{i}w_{t,i}E_{i}}$ [14]. Therefore, iterative sampling can be performed via Langevin dynamics:
@@
where $\gamma_{k}$ controls the step size and $\xi_{k}$ introduces Gaussian noise. Note that we can bridge EBM with score-matching diffusion models [37, 12, 25, 33], which updates Equ. 2 as
@@
Compared to discrete MoE routing, our compositional approach avoids routing instability and expert imbalance [22] by assigning continuous, observation-dependent weights to all components, rather than selecting a hard subset. In MoE, only a few experts are activated at each step, which can lead to underutilization of some experts and overfitting or saturation in others, especially when routing distributions are sharp or poorly calibrated. In contrast, our method aggregates contributions from all components via soft score-weighted composition, ensuring all modules remain active during optimization. Additionally, because all components participate in every training step, they receive gradient signals consistently, which encourages functional specialization.
@@ III-C Multitask Learning and Adaptation
III-C1 Multitask Learning
This factorization is particularly well-suited for multitask imitation learning, where action distributions are inherently multimodal due to diverse object properties, contact dynamics, and task goals. In contrast to monolithic policies that must capture all modes simultaneously, FDP distributes complexity evenly across diffusion components, each modeling a coherent subspace of behaviors. Unlike MoE policies, where skills may span combinations of experts across layers, our formulation yields disentangled sub-skills.
III-C2 Adapting to New Tasks
The modularity of FDP also enables efficient adaptation to unseen tasks. Instead of retraining the full model, we adapt by introducing a new diffusion component $\bm{\varepsilon}_{\theta_{\text{new}}}$, initialized via upcycling [22] – copying weights from existing components. The updated score function becomes:
@@ IV-A Experiments Setup
((b)) hang-X
((c)) cube-X
@@
We evaluate policies on 30+ tasks in simulation across MetaWorld [43], RLBench [19], and LIBERO [23]. Real-world experiments use a UR5e arm with a Robotiq gripper and a RealSense D415 camera (Fig. 2(a)). We evaluate policies on 4 distinct tasks: cube red, cube blue, hang low, and hang high. The tasks cube-X involve picking up a cube of color X from the tabletop and placing it into a designated bowl. The hang-X tasks require the robot to grasp a mug from the tabletop and precisely hang it on the X branch of a mug stand positioned on the table. Illustrations and setups of these real-world tasks are shown in Fig. 2. For MetaWorld and RLBench, goals are implicitly specified by the scene configuration, whereas for LIBERO and real-world experiments, tasks are identified using discrete integer task indices.
@@ IV-B Implementations
All policies take RGB images and joint angles as input and predict absolute joint angle trajectories. A history window of size 2 is used, with 16-step trajectories predicted and 8 steps executed. While we employ DDPM [17], FDP is solver-agnostic; alternative samplers like DDIM [32] offer comparable performance with reduced inference latency. We compare FDP against three baselines: DP [10], a monolithic diffusion policy; SDP [39], a MoE-based diffusion policy with observation-conditioned routing; and MoDE [30], a MoE variant with routing based on noise levels. We follow the original configurations used in each baselines, and proportionally reduce the model size of MoDE to match others. We refer readers to the original papers for more details on architecture and training. In FDP, four U-Net diffusion modules are composed. For adaptation, we adopt the upcycling strategy [22] to initialize new MoE experts or diffusion components from existing ones.
@@ IV-C Multitask Learning
IV-C1 Simulation
We evaluate FDP on 6 MetaWorld tasks (25 demonstrations each) and 8 RLBench tasks (50 demonstrations each). All methods are evaluated over 40 rollouts per task, with results shown in Table I. The DP baseline performs surprisingly well, particularly on tasks like drawer open, assembly, and hammer, which primarily involve reaching and grasping and exhibit fewer multimodal behaviors – making them easier to solve with a single model. Among modular baselines, SDP underperforms due to instability common in training MoE architectures [22]: too few experts limit expressiveness, while too many can cause overfitting and noisy routing. MoDE performs reasonably by routing based on the noise level, but still inherits instability from MoE training [30]. In contrast, FDP’s compositional structure avoids abrupt routing decisions by continuously composing diffusion component outputs via score-weighted aggregation, which enables stable training and more balanced component specialization of the multimodal action distributions.
Policy Cube Red Hang Low
DP 14 / 20 16 / 20
SDP 15 / 20 13 / 20
MoDE 14 / 20 16 / 20
FDP 15 / 20 17 / 20